OpenStudy (anonymous):

It takes 7 hr. for 9 cooks to prepare the food for a wedding rehearsal dinner how long will it take 2 cooks to prepare them?

OpenStudy (anonymous):

9 cooks=7hrs....so 1 cook=7/9 hr, 2 cooks=2(7/9)=14/9 hours=1 hr, 33.3 min....